# Hierarchical Decision Mamba

### v1

@@ Abstract
Recent advancements in imitation learning have been largely fueled by the integration of sequence models, which provide a structured flow of information to effectively mimic task behaviours. Currently, Decision Transformer (DT) and subsequently, the Hierarchical Decision Transformer (HDT), presented Transformer-based approaches to learn task policies. Recently, the Mamba architecture has shown to outperform Transformers across various task domains. In this work, we introduce two novel methods, Decision Mamba (DM) and Hierarchical Decision Mamba (HDM), aimed at enhancing the performance of the Transformer models. Through extensive experimentation across diverse environments such as OpenAI Gym and D4RL, leveraging varying demonstration data sets, we demonstrate the superiority of Mamba models over their Transformer counterparts in a majority of tasks. Results show that HDM outperforms other methods in most settings. The code can be found at https://github.com/meowatthemoon/HierarchicalDecisionMamba.
@@ 1 Introduction
Figure 1: HDM framework. The high-level mechanism guides the low-level controller through the task by selecting sub-goal states, based on the history of sub-goals and states. The low-level controller is conditioned on the history of past states, sub-goals, and actions to select the appropriate action. By reaching each sub-goal, the controller gets closer to completing the task.
@@
We conduct experiments on seven tasks from D4RL benchmark. Our results show that the performance of the DM is no longer dependent on the sequence of rewards, requiring no user intervention or task knowledge. Although, including the sequence of rewards does increase DM’s performance slightly. Both DM and HDM outperform their Transformer state-of-the-art predecessors.
@@
    We present the HDM and DM, the Mamba successors of HDT and DT, respectively. The models are simpler, smaller, faster and more accurate than their Transformer predecessors.
@@
    We show that DM does not require the sequence of rewards to perform, unlike its predecessor DT. This removes the need for task knowledge to craft an adequate desired reward value and user intervention. However, using the sequence of rewards does result in a slight increase in performance.
@@
    We evaluate the HDM and DM on seven environments and data sets from the D4RL [2] benchmark. Our methods outperforms the original Transformer baselines, proving the viability of the Mamba architecture for imitation learning. Recent advancements in imitation learning have been largely fueled by the integration of sequence models, which provide a structured flow of information to effectively mimic task behaviours. Currently, Decision Transformer (DT) and subsequently, the Hierarchical Decision Transformer (HDT), presented Transformer-based approaches to learn task policies. Recently, the Mamba architecture has shown to outperform Transformers across various task domains. In this work, we introduce two novel methods, Decision Mamba (DM) and Hierarchical Decision Mamba (HDM), aimed at enhancing the performance of the Transformer models. Through extensive experimentation across diverse environments such as OpenAI Gym and D4RL, leveraging varying demonstration data sets, we demonstrate the superiority of Mamba models over their Transformer counterparts in a majority of tasks. Results show that HDM outperforms other methods in most settings.
@@ 4 Methodology
We propose to replace the Transformer architecture of both the and HDT methods with the Mamba architecture, resulting in the DM and the HDM. This substitution, simplifies the overall architecture of the methods by removing the need for a causal mask and for positional encoding. Furthermore, for DM this causes the model to no longer rely on the sequence of RTG to perform the task, unlike the DT. We further explain both the DM and the HDM in the following sub-sections.
@@ 4.1 Decision Mamba
We assume to have access to a demonstration data set $D$ composed of trajectories with states, actions and rewards. Similarly to the DT, we condition the model on the sequences of states actions and RTG, each sequence with a context length of $K$. However, we perform experiments that show that the DM is also able to perform without the sequence of RTG, with a small cost in performance. During batch data sampling, a demonstration trajectory is randomly selected, followed by the choice of a trajectory index $t\in[0,T-1]$, where $T$ is the length of the trajectory. The $K-1$ states, actions and RTG preceding $t$ compose the three sequences. We also sample a right-shifted sequence of actions. This sequence corresponds to the ground-truth actions that the model should predict. The sequences are then padded with zero-filled vectors to the right, if their length is smaller than the context length. Because of this padding, we create a binary padding mask. This is a sequence of length $K$ with zeros in indexes where the previous sequences were padded. This padding mask is solely used to prevent the loss function from considering padded values, and is not used by the DM model or during inference.
@@
DM employs a Mamba architecture with repeating Mamba layers each receiving the embeddings of the previous layer and outputting new embedding vectors. We perform tests varying the number of layers, the embedding size, and the context length. There wasn’t an overall configuration that resulted in superior performance. In the experiments section, we use 6 Mamba layers, an embedding size of 128, and a context length of 20. The embeddings returned by the final Mamba layer pass through a feed-forward layer to project them to the action space of the task. Subsequently, a hyperbolic tangent activation function normalises the values between -1 and 1, which are then multiplied by the respective action range of the task.
@@ 4.2 Hierarchical Decision Mamba
The sequences of states and sub-goals each pass through a linear layer, converting them to the embedding dimension used by the model. They are then joined and passed to the Mamba layers of the high-level mechanism. The output vectors of the final Mamba layer are passed to a linear layer which projects them to the state space of the task. Then, a hyperbolic tangent activation function normalises the values between -1 and 1. Following the HDT, the states of the data set are normalised between -1 and 1. The architecture is represented in Fig. 1. The high-level mechanism is trained using the L2 loss between the predicted sub-goal sequence and the ground-truth sub-goal sequence.
@@ 5 Experiments
Table 1: Maximum accumulated returns of the DT, HDT, DM, DM with RTG, and HDM methods using 6 layers, an embedding size of 128 and context length of 20. We test the models on seven tasks from the D4RL [2] benchmark and vary the demonstration data sets.
@@
Ant expert 2715.68 2723.03 2728.36 2739.51 1853.28 2735.34 2742.3 2737.7 2741.6
medium 790.48 824.09 814.49 844.51 760.23 782.97 803.63 822.58 820.37
medium-expert 1470.33 2739.5 2751.17 811.84 1769.7 2742.88 2734.04 2733.2 2737.77
medium-replay 595.58 712.07 750.47 739.79 639.22 694.4 716.67 541.53 659.95
Antmaze large-diverse 0.0 0.0 0.1 0.0 0.0 0.0 0.0 0.0 0.0
medium-diverse 0.1 0.1 0.4 0.1 0.0 0.1 0.5 0.1 0.1
umaze 0.9 0.9 1.0 0.6 0.8 0.8 0.9 0.9 1.0
umaze-diverse 0.9 1.0 0.5 0.5 0.9 0.9 0.9 0.8 1.0
HalfCheetah expert 1470.66 1448.96 1514.67 1446.5 1025.66 1570.21 1573.29 1539.35 1601.98
medium 649.16 655.4 659.96 668.74 597.75 676.78 664.15 661.56 687.15
medium-expert 1357.27 1517.4 1519.3 895.17 1097.76 1541.82 1566.91 1595.09 1588.63
medium-replay 802.93 1071.56 1078.86 1002.38 607.06 749.42 916.15 825.28 971.4
Hopper expert 2213.17 2313.12 2294.28 2187.96 1211.95 2183.71 2218.42 1914.13 2103.62
medium 1092.03 1184.59 1185.22 1119.75 1101.02 1299.97 1289.51 1372.5 1262.85
medium-expert 2052.89 2035.33 2335.08 1733.15 1128.34 1776.79 1914.85 1569.32 1541.87
medium-replay 720.29 507.37 585.64 627.95 822.5 1065.73 1175.59 654.96 773.8
Kitchen complete 2.3 2.7 2.1 2.5 3.1 2.5 3.0 2.9 2.0
mixed 2.4 2.4 2.7 2.4 2.5 2.7 2.5 2.5 2.7
partial 2.3 2.9 2.3 3.1 2.1 3.0 3.0 2.7 3.2
Maze2d large 49.4 49.7 129.5 25.8 32.7 69.9 110.8 27.4 48.0
medium 48.8 53.1 51.0 38.9 37.8 56.3 111.9 58.6 122.2
open 12.6 16.0 17.4 18.3 25.7 30.2 32.4 13.1 17.1
umaze 47.6 51.7 93.2 42.9 58.8 86.9 186.1 48.5 24.4
Walker2d expert 1876.79 1877.02 1882.46 1879.46 1088.01 1866.81 1881.46 1853.05 1854.71
medium 1024.75 1080.22 1034.02 1074.8 932.2 1008.24 1088.64 874.03 1053.24
medium-expert 1074.54 1884.79 1879.55 1098.09 1055.61 1861.45 1866.47 1038.76 1798.05
medium-replay 1060.18 1317.15 1348.97 1256.18 716.23 1302.54 1330.53 742.24 1096.87
@@
We train each model for 1 million epochs, using batch sizes of 16, and a learning rate of $1e^{-4}$. To mitigate the influence of outliers and the inherent seed-dependency of episodes, we adopt a validation strategy wherein, every one thousand epochs, we validate the model on 100 episodes, and compute the average accumulated rewards. The results shown in the tables, represent the highest average accumulated rewards seen throughout the 1 million epochs.
@@
The results depicted in the table highlight that the DM does not require the sequence of RTG for effective performance. This is also true across different architectures as shown in Fig. 4. Notably, in the DT, eliminating this sequence impedes the DT’s ability to learn the task entirely [4]. This shows that the evolutionary parameter of the Mamba architecture successfully replaces the need for RTG. However, results also show that the DM with the sequence of RTG does seem to reach a slightly better performance than without.
Lastly, we compare the new proposed Mamba methods with their Transformer predecessors. According to the results in Table 1, the DM outperforms the DT in 15 out of the 27 settings. Additionally, the HDM outperforms the HDT in 16 out of the 27 settings. These results show that the proposed Mamba methods improve upon the Transformer predecessors in the D4RL benchmark. Overall, the HDM is the best performing model of the set. Although, the HDT and even the DT are still very competitive, it is worth noting that these models require the extra steps of causal masking and positional encoding in their implementation. Furthermore, the DT requires the specification of the value of the desired reward which is crucial to its performance.

### v2

# Decision Mamba Architectures

@@ Abstract
Recent advancements in imitation learning have been largely fueled by the integration of sequence models, which provide a structured flow of information to effectively mimic task behaviours. Currently, Decision Transformer (DT) and subsequently, the Hierarchical Decision Transformer (HDT), presented Transformer-based approaches to learn task policies. Recently, the Mamba architecture has shown to outperform Transformers across various task domains. In this work, we introduce two novel methods, Decision Mamba (DM) and Hierarchical Decision Mamba (HDM), aimed at enhancing the performance of the Transformer models. Through extensive experimentation across diverse environments such as OpenAI Gym and D4RL, leveraging varying demonstration data sets, we demonstrate the superiority of Mamba models over their Transformer counterparts in a majority of tasks. Results show that DM outperforms other methods in most settings. The code can be found at https://github.com/meowatthemoon/DecisionMamba.
@@ 1 Introduction
Figure 1: The DM architecture on the left and the HDM architecture on the right side. The DM is conditioned on the sequence of past states and actions to predict the correct action. The HDM is composed of two modules. The high-level mechanism guides the low-level controller through the task by selecting sub-goal states, based on the history of sub-goals and states. The low-level controller is conditioned on the history of past states, sub-goals, and actions to select the appropriate action.
@@
We conduct experiments on seven tasks from D4RL benchmark. Our results show that the performance of the DM is no longer dependent on the sequence of rewards, requiring no user intervention or task knowledge. Unlike the other methods, the DM does not require access to the reward function. Both DM and HDM outperform their Transformer state-of-the-art predecessors.
@@
    We present the HDM and DM, the Mamba successors of HDT and DT, respectively. The models are smaller, faster and more accurate than their Transformer predecessors.
@@
    We show that DM does not require the sequence of rewards to perform, unlike its predecessor DT. This removes the need for a reward function, task knowledge to craft an adequate desired reward value and user intervention.
@@
    We evaluate the HDM and DM on seven environments and data sets from the D4RL [2] benchmark. Our methods outperform the original Transformer baselines, proving the viability of the Mamba architecture for imitation learning. Recent advancements in imitation learning have been largely fueled by the integration of sequence models, which provide a structured flow of information to effectively mimic task behaviours. Currently, Decision Transformer (DT) and subsequently, the Hierarchical Decision Transformer (HDT), presented Transformer-based approaches to learn task policies. Recently, the Mamba architecture has shown to outperform Transformers across various task domains. In this work, we introduce two novel methods, Decision Mamba (DM) and Hierarchical Decision Mamba (HDM), aimed at enhancing the performance of the Transformer models. Through extensive experimentation across diverse environments such as OpenAI Gym and D4RL, leveraging varying demonstration data sets, we demonstrate the superiority of Mamba models over their Transformer counterparts in a majority of tasks. Results show that DM outperforms other methods in most settings.
@@ 4 Methodology
We propose to replace the Transformer architecture of both the and HDT methods with the Mamba architecture, resulting in the DM and the HDM. This substitution, simplifies the overall architecture of the methods by removing the need for a causal mask and for positional encoding. Furthermore, for DM this causes the model to no longer rely on the sequence of RTG to perform the task, unlike the DT. We further explain both the DM and the HDM in the following sub-sections. The architectures of DM and HDM are represented in Fig. 1.
@@ 4.1 Decision Mamba
We assume to have access to a demonstration data set $D$ composed of trajectories with states, actions and rewards. Similarly to the DT, we condition the model on the sequences of states actions and RTG, each sequence with a context length of $K$. However, we perform experiments that show that the DM is also able to perform without the sequence of RTG, while improving the performance. During batch data sampling, a demonstration trajectory is randomly selected, followed by the choice of a trajectory index $t\in[0,T-1]$, where $T$ is the length of the trajectory. The $K-1$ states, actions and RTG preceding $t$ compose the three sequences. We also sample a right-shifted sequence of actions. This sequence corresponds to the ground-truth actions that the model should predict. The sequences are then padded with zero-filled vectors to the right, if their length is smaller than the context length. Because of this padding, we create a binary padding mask. This is a sequence of length $K$ with zeros in indexes where the previous sequences were padded. This padding mask is solely used to prevent the loss function from considering padded values, and is not used by the DM model or during inference.
@@
DM employs a Mamba architecture with repeating Mamba layers each receiving the embeddings of the previous layer and outputting new embedding vectors. We perform tests varying the number of layers, the embedding size, and the context length. There wasn’t an overall configuration that resulted in superior performance across all the tasks. In the experiments section, we use 6 Mamba layers, an embedding size of 128, and a context length of 20. The embeddings returned by the final Mamba layer pass through a feed-forward layer to project them to the action space of the task. Subsequently, a hyperbolic tangent activation function normalises the values between -1 and 1, which are then multiplied by the respective action range of the task.
@@ 4.2 Hierarchical Decision Mamba
The sequences of states and sub-goals each pass through a linear layer, converting them to the embedding dimension used by the model. They are then joined and passed to the Mamba layers of the high-level mechanism. The output vectors of the final Mamba layer are passed to a linear layer which projects them to the state space of the task. Then, a hyperbolic tangent activation function normalises the values between -1 and 1. Following the HDT, the states of the data set are normalised between -1 and 1. The high-level mechanism is trained using the L2 loss between the predicted sub-goal sequence and the ground-truth sub-goal sequence.
@@ 5 Experiments
Table 1: Maximum accumulated returns of the DT, HDT, DM, DM with RTG, and HDM methods using 6 layers, an embedding size of 128 and context length of 20. We test the models on seven tasks from the D4RL [2] benchmark and vary the demonstration data sets. Highest values are highlighted in bold.
@@
Ant expert 2722.84 $\pm$ 9.27 2731.52 $\pm$ 3.91 2733.84 $\pm$ 9.3 2731.83 $\pm$ 8.38 1853.28 $\pm$ 7.54 2735.34 $\pm$ 11.45 2742.30 $\pm$1 14.37 2746.57 $\pm$ 10.69 2749.2 $\pm$ 11.62
medium 787.54 $\pm$ 22.87 801.41 $\pm$ 14.61 800.91 $\pm$ 17.32 787.23 $\pm$ 33.33 760.23 $\pm$ 21.73 782.97 $\pm$ 13.83 803.63 $\pm$ 14.76 821.03 $\pm$ 18.21 820.39 $\pm$ 23.41
medium-expert 1525.61 $\pm$ 724.43 2718.68 $\pm$ 11.66 2713.55 $\pm$ 15.06 2171.27 $\pm$ 792.37 1769.70 $\pm$ 18.46 2742.88 $\pm$ 13.64 2734.04 $\pm$ 15.92 2738.41 $\pm$ 8.42 2730.01 $\pm$ 12.45
medium-replay 650.23 $\pm$ 63.67 732.89 $\pm$ 13.76 815.53 $\pm$ 42.72 712.61 $\pm$ 39.5 639.22 $\pm$ 20.49 694.40 $\pm$ 13.69 716.67 $\pm$ 37.54 716.85 $\pm$ 33.65 728.2 $\pm$ 48.64
Antmaze large-diverse 0.02 $\pm$ 0.04 0.05 $\pm$ 0.05 0.05 $\pm$ 0.05 0.02 $\pm$ 0.04 0.0 $\pm$ 0.01 0.0 $\pm$ 0.01 0.0 $\pm$ 0.01 0.08 $\pm$ 0.13 0.05 $\pm$ 0.05
medium-diverse 0.2 $\pm$ 0.17 0.18 $\pm$ 0.19 0.2 $\pm$ 0.17 0.12 $\pm$ 0.04 0.0 $\pm$ 0.05 0.1 $\pm$ 0.05 0.5 $\pm$ 0.05 0.2 $\pm$ 0.17 0.15 $\pm$ 0.05
umaze 0.92 $\pm$ 0.04 0.92 $\pm$ 0.04 0.92 $\pm$ 0.04 0.52 $\pm$ 0.25 0.8 $\pm$ 0.04 0.8 $\pm$ 0.04 0.9 $\pm$ 0.04 0.95 $\pm$ 0.05 1.0 $\pm$ 0.0
umaze-diverse 0.92 $\pm$ 0.04 0.95 $\pm$ 0.05 0.9 $\pm$ 0.07 0.50 $\pm$ 0.25 0.9 $\pm$ 0.04 0.9 $\pm$ 0.07 0.9 $\pm$ 0.05 0.92 $\pm$ 0.04 1.0 $\pm$ 0.0
HalfCheetah expert 1516.19 $\pm$ 25.85 1530.86 $\pm$ 21.28 1497.99 $\pm$ 48.14 1479.88 $\pm$ 22.09 1025.66 $\pm$ 23.74 1570.21 $\pm$ 27.80 1573.29 $\pm$ 46.12 1585.85 $\pm$ 19.24 1538.32 $\pm$ 50.83
medium 655.96 $\pm$ 8.55 658.54 $\pm$ 2.02 667.08 $\pm$ 16.17 669.73 $\pm$ 15.74 597.75 $\pm$ 7.69 676.78 $\pm$ 3.14 664.15 $\pm$ 15.39 677.84 $\pm$ 13.6 679.03 $\pm$ 14.74
medium-expert 1432.58 $\pm$ 158.2 1572.42 $\pm$ 68.04 1619.32 $\pm$ 43.02 1103.51 $\pm$ 141.89 1097.76 $\pm$ 68.47 1541.82 $\pm$ 58.53 1566.91 $\pm$ 83.36 1636.29 $\pm$ 27.28 1541.19 $\pm$ 73.06
medium-replay 806.58 $\pm$ 175.57 1074.72 $\pm$ 11.14 1100.29 $\pm$ 15.2 946.68 $\pm$ 91.13 607.06 $\pm$ 110.68 749.42 $\pm$ 13.93 916.15 $\pm$ 67.71 862.53 $\pm$ 35.02 826.06 $\pm$ 96.8
Hopper expert 2129.4 $\pm$ 116.64 2234.25 $\pm$ 54.94 2239.86 $\pm$ 58.55 2072.71 $\pm$ 69.58 1211.95 $\pm$ 123.71 2183.71 $\pm$ 58.71 2218.42 $\pm$ 64.48 2260.68 $\pm$ 22.36 2230.41 $\pm$ 65.08
medium 1144.2 $\pm$ 100.69 1241.6 $\pm$ 188.58 1252.95 $\pm$ 99.69 1175.3 $\pm$ 63.39 1101.02 $\pm$ 97.47 1299.97 $\pm$ 143.74 1289.51 $\pm$ 89.36 1383.51 $\pm$ 127.34 1231.19 $\pm$ 108.38
medium-expert 2056.43 $\pm$ 166.75 2270.67 $\pm$ 62.89 2259.2 $\pm$ 117.52 1549.4 $\pm$ 112.42 1128.34 $\pm$ 128.57 1776.79 $\pm$ 86.91 1914.85 $\pm$ 105.72 2026.48 $\pm$ 180.21 1948.04 $\pm$ 342.99
medium-replay 636.21 $\pm$ 91.45 564.7 $\pm$ 99.69 648.13 $\pm$ 172.74 378.57 $\pm$ 176.15 822.5 $\pm$ 86.30 1065.73 $\pm$ 90.73 1175.59 $\pm$ 95.76 1199.42 $\pm$ 192.37 770.51 $\pm$ 95.49
Kitchen complete 2.52 $\pm$ 0.18 2.53 $\pm$ 0.13 2.42 $\pm$ 0.38 2.7 $\pm$ 0.23 3.1 +0.16 2.5 $\pm$ 0.09 3.0 $\pm$ 0.02 2.58 $\pm$ 0.36 2.2 $\pm$ 0.14
mixed 2.28 $\pm$ 0.27 2.55 $\pm$ 0.21 2.15 $\pm$ 0.34 2.28 $\pm$ 0.13 2.5 $\pm$ 0.26 2.7 $\pm$ 0.18 2.5 $\pm$ 0.37 2.8 $\pm$ 0.2 2.65 $\pm$ 0.09
partial 2.65 $\pm$ 0.27 3.0 $\pm$ 0.07 2.08 $\pm$ 0.51 2.42 $\pm$ 0.44 2.1 $\pm$ 0.25 3.0 $\pm$ 0.03 3.0 $\pm$ 0.02 3.02 $\pm$ 0.04 2.55 $\pm$ 0.43
Maze2D large 106.72 $\pm$ 17.23 103.7 $\pm$ 15.03 103.45 $\pm$ 19.85 34.1 $\pm$ 8.97 32.7 $\pm$ 16.05 69.9 $\pm$ 14.08 110.8 $\pm$ 17.74 79.65 $\pm$ 20.43 93.7 $\pm$ 71.77
medium 40.07 $\pm$ 9.19 33.72 $\pm$ 13.3 154.0 $\pm$ 63.45 33.72 $\pm$ 7.87 37.8 $\pm$ 5.34 56.3 $\pm$ 5.35 111.9 $\pm$ 45.34 114.1 $\pm$ 75.94 145.23 $\pm$ 20.81
open 16.62 $\pm$ 0.48 17.17 $\pm$ 1.49 15.85 $\pm$ 3.61 19.08 $\pm$ 2.17 25.7 $\pm$ 1.64 30.2 $\pm$ 8.63 32.4 $\pm$ 4.56 32.33 $\pm$ 2.05 26.67 $\pm$ 0.76
umaze 58.68 $\pm$ 20.14 59.3 $\pm$ 20.61 61.2 $\pm$ 19.74 37.72 $\pm$ 11.7 58.8 $\pm$ 17.03 86.9 $\pm$ 10.47 186.1 $\pm$ 16.43 110.25 $\pm$ 51.04 31.13 $\pm$ 2.58
Walker2D expert 1863.25 $\pm$ 11.68 1867.61 $\pm$ 9.43 1877.92 $\pm$ 5.12 1869.07 $\pm$ 8.54 1088.01 $\pm$ 14.72 1866.81 $\pm$ 4.72 1881.46 $\pm$ 9.46 1874.51 $\pm$ 10.86 1861.86 $\pm$ 7.2
medium 1070.96 $\pm$ 61.03 1081.54 $\pm$ 42.68 1175.83 $\pm$ 58.49 1106.63 $\pm$ 24.09 932.2 $\pm$ 40.81 1008.24 $\pm$ 54.91 1088.64 $\pm$ 43.09 $1$.94 $\pm$ 32.08 1093.42 $\pm$ 36.83
medium-expert 1310.26 $\pm$ 352.94 1883.23 $\pm$ 19.77 1880.87 $\pm$ 24.98 1153.5 $\pm$ 54.57 1055.61 $\pm$ 35.89 1861.45 $\pm$ 23.23 1866.47 $\pm$ 18.67 1871.34 $\pm$ 4.88 1843.48 $\pm$ 11.49
medium-replay 1129.27 $\pm$ 129.82 1335.64 $\pm$ 10.87 1377.9 $\pm$ 21.19 1197.66 $\pm$ 51.96 716.23 $\pm$ 78.38 1302.54 $\pm$ 12.94 1330.53 $\pm$ 34.71 1310.88 $\pm$ 32.54 1237.97 $\pm$ 63.21
@@
We train each model for 1 million epochs, using batch sizes of 16, and a learning rate of $1e^{-4}$. To mitigate the influence of outliers and the inherent seed-dependency of episodes, we adopt a validation strategy wherein, every one thousand epochs, we validate the model on 100 episodes, and compute the average accumulated rewards. Due to the seed dependency, we repeat each experiment across 4 different random seeds. The presented results are the average values across the 4 seeds of the highest accumulated rewards seen throughout the 1 million epochs.
@@
The results depicted in the table highlight that the DM does not require the sequence of RTG for effective performance. Additionally, results also show that the DM without the sequence of RTG does seem to reach a slightly better performance than with the sequence. This is also true across different architectures as shown in Fig. 4. Notably, in the DT, eliminating this sequence impedes the DT’s ability to learn the task entirely as shown in [4]. This indicates that the evolutionary parameter of the Mamba architecture successfully replaces the need for RTG. Since DM without the sequence of rewards achieves higher performance without requiring additional user interaction, we can conclude that DM should be used without the reward sequence.
Lastly, we compare the new proposed Mamba methods with their Transformer predecessors. According to the results in Table 1, the DM without rewards outperforms the DT in 15 out of the 27 settings. Additionally, the HDM outperforms the HDT in 22 out of the 27 settings. The superiority of the Mamba models exists even while comparing to the DT with the ideal desired reward for each task. When comparing the DM to the DT using a fixed desired reward of 10k, the DM outperforms the DT in 17 out of the 27 settings. These results show that the proposed Mamba methods improve upon the Transformer predecessors in the D4RL benchmark. Overall, the DM without rewards is the best performing model of the set. Although, the HDM and the DT are still very competitive, it is worth noting that HDM requires two models and pre-processing the data set, while DT requires user interaction and task knowledge. Moreover, unlike the other models, DM can be applied to tasks without a reward function.